A ball is thrown vertically upwards with a velocity of 10m/s . It returns to the ground with the velocity of 9m/s . If g=9.8m/s^2 ,then the max height attained by the ball is nearly?

Grade:10

2 Answers

Ram Prakash
5 years ago
u=10 m/s
a =-9.8 m/s^2( since g acts downwards)
v= 0 
since 
v^2 = u^2 +2as
therefore
s= v^2 -u^2 / 2 a
  = 0- 100 /2×-9.8
  = -100 / -19.6 
 = 5.1020 m 
  = -100 / 
 
Khimraj
5 years ago

Let the mass of the body be M. It is thrown up with a velocity of 10 m/s, but returns to the ground with a velocity of 9 m/s.

It means that in upward and downward journey it lost energy due to dissipative forces equal to kinetic energy of 1m/s, i.e. it lost energy equivalent to kinetic energy of ½ m/s either way.

So it is equivalent to saying that it was thrown up with an initial velocity of 9.5 m/s, rather than 10m/s, and forget about dissipative forces.

At the highest point the entire kinetic energy is converted into potential energy. Let the height achieved be h. Then

½ m v² = m g h,==> h = v²/2g = 9.5²/2×9.8= 4.61 m.

The body shall achieve a height of 4.61 m.